What is 34m3 in micro m3?
1
Expert's answer
2019-08-20T10:25:16-0400
"1\mu m=10^{-6}m"

"34m^3=34(10^6\mu m)^3=34\cdot 10^{18}\mu m^3"

"34m^3=3.4\cdot 10^{19}\mu m^3"